# Search for a gravitational wave background from primordial black hole binaries using data from the first three LIGO-Virgo-KAGRA observing runs

### v2

@@ Abstract
Using the cross-correlation data from the first three observing runs of the LIGO-Virgo-KAGRA Collaboration, we search for a gravitational-wave background (GWB) from primordial black holes, arising from the superposition of compact binary coalescence events. We consider both early and late binary formation mechanisms, and perform Bayesian parameter inference, investigating different prior distributions of the model parameters. From the non-detection of the GWB, we provide constraints on the fraction of primordial black holes contributing to the present dark matter energy density.
@@
Another approach is to search unresolvable binaries as a GWB. Some studies have explored constraints on PBHs by considering the GWB produced by their mergers, a line of research pioneered in [51]. Constraints have been discussed in the literature [26, 28, 52, 53, 54, 55, 56, 57], typically involving a rough comparison between theoretical predictions and the upper bounds obtained from the isotropic GWB search performed by the LVK collaboration [58, 59, 60], which assumes a power-law spectrum $\Omega_{\rm GW}\propto f^{\alpha}$. Although this assumption may be valid for a GWB originating from an unresolved BBH population from Population I or II stars (where lower end is bounded by the Chandrasekhar limit, while the intermediate mass range is constrained by pair-instability), PBH mass distributions can extend significantly beyond $\sim 50\,M_{\odot}$. As a result, the GWB spectrum could deviate from a simple power-law form when PBH sources are considered ([57]).
Motivated by these, we analyze public GWB data from the first three LVK observing runs (O1–O3) [58, 59, 60] and perform a full Bayesian parameter inference for the GWB from PBHs with a wide range of masses, including sub-solar masses up to $10^{3}$ solar masses. This allows us to account for the complete spectral shape and explore a broad parameter space that extends beyond the detection of individual mergers by LVK. Some recent works in this direction are [43, 57], although the former was primarily focused on distinguishing between astrophysical and primordial BBHs. In addition, the PBH model used in [43, 57] were limited to a single mass bin of $1$ and $30M_{\odot}$. It is worth noting that in what follows it is for the first time considered the combination of different formation channels to get corresponding constraints and assess their relative significance.
@@ III.2 Prior Distributions
Figure 1: (Left) 95% C.L. constraints on $f_{\rm PBH}$ as a function of $M_{\rm PBH}$($=\mu$) from the search for the GWB using the different $R_{\rm clust}$ choices are shown in red ($1$-dashed, $4\times 10^{2}$-solid, $10^{3}$-dotted). We also display supernova lensing constraints (SNe) [94] in purple, Massive Compact Halo Object (MACHO) constraints [95] in orange designated as (EROS), the recent Optical Gravitational Lensing Experiment (OGLE) constraints are shown in green with different Milky Way (MW) models [96], and cosmic microwave background (CMB) constraints from [97]. We used the PBHBounds [98] to generate this plot. (Right) Zoomed‑in view of the same plot.
@@ III.3 Result
The results indicate that no significant GWB associated with PBHs or CBCs has been detected over the noise hypothesis, with logarithmic Bayes factors $\sim-1$. We establish upper limits on the CBC energy density parameter, $\Omega_{\rm cbc}$, of $\sim 3\times 10^{-9}$ at the $95\%$ C.L. level which are consistent with LVK constraints [60].
From the non-detection, we can constrain the fraction of PBHs, $f_{\rm PBH}$, as a function of the mean mass $\mu$. The obtained constraint, along with other studies from lensing surveys like EROS ([95]), OGLE ([96]), supernova Ia ([94]) and cosmic microwave background (CMB) constraints for different $R_{\rm clust}$ choices are summarized in Fig. 1 and II. We also provide the resulting corner plots for all model parameters in Appendix D.
@@
$1$ $3.6\times 10^{-2}$ $5.3\times 10^{-3}$ $1.0\times 10^{0}$
@@ Acknowledgments
where $\sigma^{2}_{M}=0.005$ is the variance of matter density perturbations, evaluated at binary formation, and $\bar{N}$ is the number of PBHs within a sphere of comoving radius around the initial PBH pair, given by
@@
$\hat{\sigma}^{IJ}_{t_{i},f_{j}}=\frac{1}{2T\Delta f}\cdot\frac{P_{1,t_{i}}(f_{j})\,P_{2,t_{i}}(f_{j})}{\gamma^{2}_{IJ}(f_{j})\,S^{2}_{0}(f_{j})},$ (17)

### v3

@@ Abstract
Using the cross-correlation data from the first three observing runs of the LIGO-Virgo-KAGRA Collaboration, we search for a gravitational-wave background (GWB) from primordial black holes, arising from the superposition of compact binary coalescence events. We consider both early and late binary formation mechanisms and perform Bayesian parameter inference. From the non-detection of the GWB, we provide constraints on the fraction of primordial black holes contributing to the present dark matter energy density.
@@
Another approach is to search unresolvable binaries as a GWB. Some studies have explored constraints on PBHs by considering the GWB produced by their mergers, a line of research pioneered in [50]. Constraints have been discussed in the literature [25, 27, 51, 52, 53, 54, 55], typically involving a rough comparison between theoretical predictions and the upper bounds obtained from the isotropic GWB search performed by the LVK collaboration [56, 57, 58], which assumes a power-law spectrum $\Omega_{\rm GW}\propto f^{\alpha}$. Although this assumption may be valid for a GWB originating from an unresolved BBH population from Population I or II stars (where lower end is bounded by the Chandrasekhar limit, while the intermediate mass range is constrained by pair-instability), PBH mass distributions can extend significantly beyond $\sim 50\,M_{\odot}$. As a result, the GWB spectrum could deviate from a simple power-law form when PBH sources are considered ([55]).
Motivated by these, we analyze public GWB data from the first three LVK observing runs (O1–O3) [56, 57, 58] and perform a full Bayesian parameter inference for the GWB from PBHs with a wide range of masses, including sub-solar masses up to $10^{3}$ solar masses. This allows us to account for the complete spectral shape and explore a broad parameter space that extends beyond the detection of individual mergers by LVK. Some recent works in this direction are [42, 55], although the former was primarily focused on distinguishing between astrophysical and primordial BBHs. In addition, the PBH model used in [42, 55] were limited to a single mass bin of 111 and $30M_{\odot}$. It is worth noting that in what follows it is for the first time considered the combination of different formation channels to get corresponding constraints and assess their relative significance.
@@ III.2 Prior Distributions
Figure 1: (Left) 95% C.L. constraints on $f_{\rm PBH}$ as a function of $M_{\rm PBH}$($=\mu$) from the search for the GWB using the different $R_{\rm clust}$ choices are shown in red (111-dashed, $4\times 10^{2}$-solid, $10^{3}$-dotted). We also display supernova lensing constraints (SNe) [92] in purple, Massive Compact Halo Object (MACHO) constraints [93] in orange designated as (EROS), the recent Optical Gravitational Lensing Experiment (OGLE) constraints are shown in green with different Milky Way (MW) models [94], and cosmic microwave background (CMB) constraints from [95]. We used the PBHBounds [96] to generate this plot. (Right) Zoomed‑in view of the same plot.
@@ III.3 Result
[…]
From the non-detection, we can constrain the fraction of PBHs, $f_{\rm PBH}$, as a function of the mean mass $\mu$. The obtained constraint, along with other studies from lensing surveys like EROS ([93]), OGLE ([94]), supernova Ia ([92]) and cosmic microwave background (CMB) constraints for different $R_{\rm clust}$ choices are summarized in Fig. 1 and Table II. We also provide the resulting corner plots for all model parameters in Appendix D.
@@
111 $3.6\times 10^{-2}$ $5.3\times 10^{-3}$ $1.0\times 10^{0}$
@@ Acknowledgments
where $\sigma_{M}=0.005$ is the variance of matter density perturbations, evaluated at binary formation, and $\bar{N}$ is the number of PBHs within a sphere of comoving radius around the initial PBH pair, given by
@@
$(\hat{\sigma}^{IJ}_{t_{i},f_{j}})^{2}=\frac{1}{2T\Delta f}\cdot\frac{P_{1,t_{i}}(f_{j})\,P_{2,t_{i}}(f_{j})}{\gamma^{2}_{IJ}(f_{j})\,S^{2}_{0}(f_{j})},$ (17)